OpenStudy (lxelle):

In a chemical reaction a compound X is formed from a compound Y. The masses in grams of X and Y present at time t seconds after the start of the reaction are x and y respectively. The sum of the two masses is equal to 100 grams throughout the reaction. At any time, the rate of formation of X is proportional to the mass of Y at that time. When t = 0, x = 5 and dx/dt = 1.9. (i) Show that x satisfies the differential equation: dx/dt = 0.02(100 - x).

OpenStudy (unklerhaukus):

\[X+Y = 100\] \[dX/dt \propto Y\\ \qquad\quad=kY\]

OpenStudy (lxelle):

I've got dx/dt= k(100-x) 1.9=k(95) But k isn't 0.02

OpenStudy (unklerhaukus):

k = 1.9/95

OpenStudy (unklerhaukus):

Why do you say that `k isn't 0.02` ?

OpenStudy (lxelle):

Oh lol sorry lol. My wrong.

OpenStudy (unklerhaukus):

all sorted then?